OpenStudy (anonymous):

solve for x and y by substitution: x/3+y/4=2 2x/3-y/2=0

OpenStudy (anonymous):

SOMEONE PLEASE HELP ! =(

OpenStudy (anonymous):

working on it

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

simplify the first one then the second

OpenStudy (anonymous):

the first one comes out to be 4x-3y=24

OpenStudy (anonymous):

the second comes out to be 4x-3y=0

OpenStudy (anonymous):

where on earth do you find the 24 from? :S

OpenStudy (anonymous):

it becomes 4x/12 + 3y/12 = 2

OpenStudy (anonymous):

its the lcm of 3 and 4 and then multiplied by 2 on right hand side

OpenStudy (anonymous):

then simplifies to 4x+3y=24

OpenStudy (anonymous):

4x-3y=2 -4x-3y=0

OpenStudy (anonymous):

same thing as 4x-3y=24 -4x+3y=0